The verbal part of the Graduate Record Exam (GRE) has a mean of 500 and a standard deviation of 100. Use the normal distribution to answer the following questions:
a. If you wanted to select only students at or above the 90th percentile, what verbal GRE score would you use as a cutoff score?
b. What verbal GRE score corresponds to a percentile rank of 15%?
c. What’s the percentile rank of a GRE score of 628?

Answer by ewatrrr(24785)

a. z = invNorm(.90) = (X-500)/100
100invNorm(.90) + 500 = X, score would you use as a cutoff score
b.
100invNorm(.15) + 500 = X, verbal GRE score corresponds to a percentile rank of 15%
z = (628-500)/100 = 128/100 = 1.28
normalcdf(-100, 1.28)